# An Efficient Finite Difference Approximation via a Double Sample-Recycling Approach

### v1

@@ B.2 Proof of Theorem 4
where the second equality is due to (5). As for the third equality, we omit the lengthy discussion on the technical assumptions needed to ensure $\mathbb{E}\left[o\left(\widehat{h}_{n}^{2}\right)\right]=o\left(n^{-1/3}\right)$5. In the following, we keep assuming that the operators $\mathbb{E}[\cdot]$ (or $\mathrm{Var}[\cdot]$) and $o(\cdot)$ (or $o_{p}(\cdot)$) can be exchanged.
@@
The third term on the RHS of (B.2)
$=$ $\frac{K^{2}n_{b}^{2}}{n^{2}}\mathrm{Var}\left[\widehat{\alpha}^{\prime}(\theta_{0})+\widehat{B}\widehat{h}_{n}^{2}\right]$
@@ B.4 Proof of Proposition 5
It is obvious that (36) holds when $n_{b}=o(n)$. In the following, we consider the bias and variance terms of case 2 (i.e., $\underset{n_{b}\to\infty}{\lim}\underset{n\to\infty}{\lim}Kn_{b}/n=r$, where $0<r\leq 1$) in Proposition 5 separately.
@@
      $-1-{\rm{sign}}(B)\leq\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}D{\bm{c^{\top}Pc^{4}}}\leq 1-{\rm{sign}}(B),$
@@
      $=$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{2/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}\left(\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}+{\rm{sign}}(B)\sigma(\theta_{0})\right)$
      $=$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{2/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}\sqrt{\frac{r}{K}}D{\bm{c^{\top}Pc^{4}}}\left(\sqrt{\frac{r}{K}}D{\bm{c^{\top}Pc^{4}}}+{\rm{sign}}(B)\sigma(\theta_{0})\right)$
@@
      $\left|\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}-\alpha^{\prime}(\theta_{0})\right|=$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{n_{b}}{n}}D{\bm{c^{\top}Pc^{4}}}n^{-1/3}+o\left(n^{-1/3}\right)$
      $\leq$ $\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/6}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{r}{K}}\left|D{\bm{c^{\top}Pc^{4}}}\right|n^{-1/3}+o\left(n^{-1/3}\right)$
      $\leq$ $\left(1+\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}\left|D{\bm{c^{\top}Pc^{4}}}\right|\right)\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/6}+o\left(n^{-1/3}\right).$
@@
B.5 A similar conclusion to Theorem 4 for explaining the variance reduction and bias reduction in Remark 2
To illustrate that variance reduction and possible bias reduction occur when substituting $\mathbb{E}[\Delta(h_{k})],k=1,...,K$ with their estimators, we refrain from using the substitution and instead assume that $\mathbb{E}\Delta(h_{k}),k=1,...,K$ are all precisely known. Then, the new DSR-CFD estimator, denoted by $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$, is also formulated as (32) but with $\Delta_{j}^{t}(h_{k})$ replaced by
$\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\mathbb{E}\Delta(h_{k})\right)+\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}.$
In the following, we provide the asymptotic properties of $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$.
@@
Proof
We prove (99) and (100) separately.
Proof of (99): It follows that
According to Appendix B.2, to prove (99), it suffices to prove the expectation of the second term on the RHS of (101) is $o\left(n^{-1/3}\right)$. The proof is as follows.
$\mathbb{E}\left[\frac{n_{b}}{n}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})\right)\right]$
$=$ $\frac{n_{b}}{n}\sum_{k=1}^{K}|h_{k}|\mathbb{E}\left[\frac{\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})}{\left|\widehat{h}_{n}\right|}\right]$
@@
The proof is complete.
Appendix C C Additional Experiment Results
Table 4: Comparison of the bias of the DSR-CFD and Opt-CFD estimators in Example 2 when $x=0.4$.
[…]
Budget $=10^{6}$ $-0.0090$ $-0.0063$ 1.4286
Budget $=2\times 10^{6}$ $-0.0067$ $-0.0051$ 1.3137
Budget $=4\times 10^{6}$ $-0.0051$ $-0.0039$ 1.3077
Budget $=8\times 10^{6}$ $-0.0040$ $-0.0031$ 1.2903

### v4

@@ B.2 Proof of Theorem 4
where the second equality is due to (5). As for the third equality, we omit the lengthy discussion on the technical assumptions needed to ensure $\mathbb{E}\left[o\left(\widehat{h}_{n}^{2}\right)\right]=o\left(n^{-1/3}\right)$LABEL:fn:uniformaly_integrable. In the following, we keep assuming that the operators $\mathbb{E}[\cdot]$ (or $\mathrm{Var}[\cdot]$) and $o(\cdot)$ (or $o_{p}(\cdot)$) can be exchanged.
@@
[…]
@@ B.4 Proof of Proposition 5
It is obvious that (36) holds when $n_{b}=o(n)$. In the following, we consider the bias and variance terms of cases 2 and 3 (i.e., $Kn_{b}/n\to r$, where $0<r\leq 1$) in Proposition 5 separately.
@@
      $-1-{\rm{sign}}(B)\leq\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}D\Lambda\leq 1-{\rm{sign}}(B),$
@@
      $=$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{2/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}\sqrt{\frac{n_{b}}{n}}D\Lambda\left(\sqrt{\frac{n_{b}}{n}}D\Lambda+{\rm{sign}}(B)\sigma(\theta_{0})\right)$
      $=$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{2/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/3}\sqrt{\frac{r}{K}}D\Lambda\left(\sqrt{\frac{r}{K}}D\Lambda+{\rm{sign}}(B)\sigma(\theta_{0})\right)$
@@
      $\left|\mathbb{E}\widehat{\Delta}_{n,\widehat{h}_{n}}-\alpha^{\prime}(\theta_{0})\right|=$ $\left(\frac{B\sigma^{2}(\theta_{0})}{4n}\right)^{1/3}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{n_{b}}{n}}D\Lambda n^{-1/3}+o\left(n^{-1/3}\right)$
      $\leq$ $\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/6}+\left(\frac{4B^{2}}{\sigma^{2}(\theta_{0})}\right)^{1/6}\sqrt{\frac{r}{K}}\left|D\Lambda\right|n^{-1/3}+o\left(n^{-1/3}\right)$
      $\leq$ $\left(1+\sqrt{\dfrac{r}{K}}\dfrac{2}{\sigma(\theta_{0})}\left|D\Lambda\right|\right)\left(\dfrac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/6}+o\left(n^{-1/3}\right).$
@@
B.5 The Bias and Variance of $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$
In this section, we provide the asymptotic properties of $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$ as follows.
@@ B.5 The Bias and Variance of $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$
Proof.
Recall that
$\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+n_{b}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})\right)+Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right),$ (105)
and we prove (103) and (104) separately.
Proof of (103): According to Appendix B.2, to prove (103), it suffices to prove the expectation of the second term on the RHS of (105) is $o\left(n^{-1/3}\right)$. The proof is as follows.
$\mathbb{E}\left[\frac{n_{b}}{n}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})\right)\right]=\frac{n_{b}}{n}\sum_{k=1}^{K}|h_{k}|\mathbb{E}\left[\frac{\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})}{\left|\widehat{h}_{n}\right|}\right]$
@@
The proof is complete. ∎
Appendix C C Additional Experiment Results in Section 6
C.1 Additional Results for Example 1
In Figure 5, we depict the sensitivity of different methods w.r.t. $r$, where the suffixes “1” and “2” denote cases 1 and 2, respectively, in Example 1. We compare the performance of Cor-CFD, EM-CFD and BOOT-CFD (bootstrap-based CFD, i.e., the estimator $\widehat{\Delta}_{n_{2},\widehat{h}_{n_{2}}}$ discussed at the beginning of Section 4). We set the number of sample pairs to 1000. When examining the EM-CFD method, we generate $h_{1},\ldots,h_{1000r}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,1\times(1000r)^{-1/5}\right)$.
As observed in Figure 5, the presence of heteroskedasticity complicates the estimation problem for each method, resulting in inferior estimations. However, the impact of heteroskedasticity is less pronounced in the Cor-CFD method, which consistently outperforms other methods across various scenarios. The EM-CFD method, on the other hand, exhibits less stability when $r$ is small, possibly due to less accurate perturbation estimation. Additionally, the Cor-CFD method demonstrates stability w.r.t. $r$, while the performance of EM-CFD and BOOT-CFD deteriorate as $r$ increases. Even when $r=1$, the Cor-CFD method maintains strong performance, making it particularly suitable for small sample sizes, where all samples are utilized for both constant and gradient estimation, effectively leveraging the available sample information.
Figure 5: Illustration of the MSE of different methods w.r.t. $r$ for cases 1 and 2 in Example 1.
C.2 Additional Results for Example 2
We compare the estimated MSEs of our proposed Cor-CFD estimator, the EM-CFD estimator and the Opt-CFD estimator in Example 2 by setting $\theta_{0}=0$ and varying the sample pairs from $n=50$ to $n=500$. When considering the EM-CFD algorithm, we set $r=0.1$ and generate $h_{1},...,h_{rn}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,(rn)^{-1/5}\right)$. When applying the Cor-CFD algorithm, we set $K=10$ and $r=1$.
Figure 6 shows the comparison results. The figure illustrates that the estimated MSEs of the three estimators decrease gradually as the sample size increases. However, compared to the Opt-CFD and Cor-CFD estimators, the EM-CFD estimator exhibits a larger MSE. Moreover, the MSE of the Cor-CFD estimator is consistently close to that of the Opt-CFD estimator. For example, when the sample-pair size is 50, the MSE of the EM-CFD estimator is approximately 3.878, while the MSEs of the Cor-CFD and Opt-CFD estimators are approximately 0.284 and 0.160, respectively. When the sample-pair size is 500, the three MSEs are 0.286, 0.045 and 0.035, respectively.
Appendix D D Simulation Settings and Additional Experiment Results in Section 7
D.1 Simulation Settings in Section 7
The simulation settings of different methods in Section 7 are detailed as follows:
  1. 1.
    The CorCFD-L-BFGS algorithm. We set the perturbation numbers $K=5$, the initial batch size $T_{0}=20$, the proportion of the pilot samples $r=1$, the bootstrap numbers $I=100$, the line search parameters $(l_{1},l_{2})=(10^{-4},0.5)$, the initial step length $a_{0}=1$ and the initial perturbation generator $\mathcal{P}_{0}$ is identical to that in Example 2.
  2. 2.
    The TraCFD-L-BFGS algorithm. We set $B=1$ and $\sigma(\theta)=1$ for any $\theta$ in the feasible domain, the line search parameters $(l_{1},l_{2})=(10^{-4},0.5)$ and the initial step length $a_{0}=1$.
  3. 3.
    The NEWUOA. Similar to the operations in Shi et al. (2023), we utilize the default parameters in R package “minqa” and repeat the NEWUOA until exhausting the allocated budget.
D.2 Discussion on the Comparison with TraCFD-L-BFGS
Figure 7: Comparison of the iteration process for $d=10$.
Figure 7 illustrates the iteration processes of the TraCFD-L-BFGS and CorCFD-L-BFGS algorithms in optimizing the Zahkarov function during a specific replication for $d=10$ and $n=10^{4}$. The TraCFD-L-BFGS algorithm use only one sample pair in each iteration, resulting in a significantly higher number of iterations compared to the CorCFD-L-BFGS algorithm. As shown in the figure, initially, the TraCFD-L-BFGS algorithm’s function value decreases rapidly, while the CorCFD-L-BFGS algorithm’s function value decreases more gradually. This is because, at the beginning, when the initial value is far from the optimum, gradient descent is more critical than precise gradient estimation. However, as iterations progress, the function value of the TraCFD-L-BFGS algorithm starts to fluctuate, whereas that of the CorCFD-L-BFGS algorithm steadily converges. Near the optimum, the true gradient value is small, making noise more significant. At this stage, obtaining an accurate gradient estimate becomes crucial for the algorithm to find the correct descent direction. Therefore, using more samples to reduce variance is more effective.
D.3 Discussion on the Comparison with the SPSA Algorithm
We further evaluate the performance of the SPSA algorithm (Spall, 1992) using the function “minimizeSPSA” in Python library “noisyopt”. The SPSA algorithm updates in the $k$th iteration as follows: $x_{k+1}=x_{k}+a_{k}[f(x_{k}+c_{k}\xi_{k})-f(x_{k}-c_{k}\xi_{k})]/c_{k}$, where $\xi_{k}$ denotes a random direction, $a_{k}=a(0.01*niter+k+1)^{-\alpha_{0}}$, $c_{k}=c(k+1)^{-\gamma_{0}}$ and $niter$ is the number of total iterations. In minimizeSPSA, $\alpha_{0}=0.602$ and $\gamma_{0}=0.101$, as suggested in (Spall, 1992). In our experiment, $niter$ is the number of sample pairs. We manually tune the hyperparameters $a$ and $c$.
To strike a balance between dimensionality and the complexity of hyperparameter tuning, we focus on the case where $d=10$. We explore 20 different combinations of $a$ and $c$ with $a=\{10^{-7},5\times 10^{-7},10^{-6},5\times 10^{-6},10^{-5}\}$ and $c=\{10^{-2},10^{-1},1,10\}$. Our findings indicate that $a=10^{-6}$ and $c=1$ yield the best results among all combinations. For sample-pair sizes of $10^{3}\times d$, $10^{4}\times d$ and $10^{5}\times d$, we observe the SGs of 1.513, 1.458 and 1.421, respectively, and the OGs of 6.581, 2.278 and 2.024, respectively. Comparing these results with those presented in Table 4, we surprisingly find that our proposed algorithm significantly outperforms the SPSA algorithm. We posit several potential explanations for this observation:
  1. 1.
    Our algorithm stems from the L-BFGS, which is a second-order algorithm, while the SPSA algorithm corresponds to a first-order algorithm. Typically, the second-order algorithm performs better than first-order one.
  2. 2.
    Another reason is the same as that discussed in Section D.2. That is, the gradient approximated by one sample pair is less accurate than that approximated by a batch of sample pairs.
It is important to note that the SPSA algorithm was originally proposed as a sample-efficient method for multi-dimensional optimization problems. However, our comparative analysis reveals an intriguing area for further investigation: the relative efficacy of using a single sample pair (as in the SPSA algorithm) versus multiple sample pairs (as in our CorCFD-L-BFGS algorithm) for gradient approximation. This discrepancy in performance suggests a valuable avenue for future research, which could potentially yield significant insights into the strengths and limitations of these approaches across diverse optimization scenarios. Such investigations may contribute to a more nuanced understanding of when and how to apply these methods most effectively, particularly in high-dimensional optimization problems.